Three infinite straight wires are fixed in place and aligned parallel to the z-axis as shown. The wire at (x.y) (-16 cm, 0d carries current 1 3 A in the negative z-4 direction. The wire at (xy) (16 cm, 0) carries current 12 0.9 A in the positive z-direction. The wire at (xy) (0, 27.7 cm) carries current 137.6 A in the positive z-direction. What is Bx(0,0), the x-component of the magnetic field produced by these three wires at the origin? You currently have 0 submissions for this question. Only 10 submission are allowed You can make 10 more submissions for this question. What is By(0,0), the y-component of the magnetic field produced by these three wires at the origin? You currently have 0 submissions for this question. Only 10 submission are allowed. You can make 10 more submissions for this question. What is Fx(1), the x-component of the force exerted on a one meter length of the wire carrying current 1? You currently have 0 submissions for this question. Only 10 submission are allowed You can make 10 more submissions for this question. What is Fy(1), the y-component of the force exerted on a one meter length of the wire carrying current l1? N Semn You currently have 0 submissions for this question. Only 10 submission are allowed You can make 10 more submissions for this question. owhat is Fx(2), the x-component of the force exerted on a one meter length of the wire carrying current l2? You currently have 0 submissions for this question. Only 10 submission are allowed. You can make 10 more submissions for this question.
Answer #1

A)

Bx(0, 0) = (u0/2 pi) (I3/0.277) = (2 x 10^-7) x (7.6/0.277) = 5.49 x 10^-6 T

B)

By(0, 0) = - (2 x 2 x 10^-7) [(0.9/0.32) + (3.8/0.32)] = - 5.88 x 10^-6 T

C)

Fx(1) = - (2 x 10^-7) (3/0.32) (0.9 + 7.6) = - 1.59 x 10^-5 N

D)

Fy(1) = (-2 x 10^-7) x 7.6 x 3 x sin60) /(2 x 0.32) = - 6.17 x 10^-6 N

E)

Fy(2) = (2 x 10^-7)(0.9/0.32)(7.6/2 - 3) = 4.5 x 10^-7 N

Comment in case any doubt please rate my answer ....